each pair of points lie on a line with the given slope. find x
(2,4), (x,8) slope=-2

Answer by tutorcecilia(2144)

(2,4), (x,8) slope=-2
(y-y2)=m(x-x2)[Use the point-slope form of a line to solve]
(y-4)=(-2)(x-2) [Plug-in the values]
y-4=-2x+4 [Solve for y]
y=-2x+4+4
y=-2x+8
.
y=-2x+8[Plug-in the value of y=8 and solve for x]
8=-2x+8
8-8=-2x+8-8
0=-2x
x=0
Point (0, 8)
.
Check by plugging the point (2,4)into the equation of the line
y=-2x+8
4=(-2)(2)+8
4=-4+8
4=4 [Checks out]